OpenStudy (anonymous):

X is a normally distributed random variable with a standard deviation of 4.00. Find the mean of X if 12.71% of the area under the distribution curve lies to the right of 13.56

OpenStudy (anonymous):

\[ P{ x > 13.56] = .1271 P[ X < 13.56] = 1- .1271=0.8729 \]

OpenStudy (anonymous):

12.3 10.3 9.0 8.5 these are my choices

OpenStudy (anonymous):

\[ P{ x > 13.56] = .1271\\ P[ X < 13.56] = 1- .1271=0.8729\\ \]

OpenStudy (anonymous):

(13.56 - mean)/4 = 1.15 mean =8.96 So 9 is the answer

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw

OpenStudy (kropot72):

@eliassaab The population mean cannot be a lower value than 13.56. Therefore none of the proposed choices is correct.

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

@kropot72 \[ \int_{13.56}^{\infty } \frac{e^{-\frac{1}{32} (x-9)^2}}{4 \sqrt{2 \pi }} \, dx=0.12714 \] Try it on any numerical integration system.

OpenStudy (anonymous):

Copy and paste on your browser http://www.wolframalpha.com/input/?i=Integrate [E^%28-%281%2F32%29+%28-9+%2B+x%29^2%29%2F%28++4+Sqrt[2+\[Pi]]%29%2C+{x%2C+13.56%2C+\[Infinity]}]

OpenStudy (anonymous):

@kropot72, may be you did not notice that the problem stated the area to the right of 13.56

OpenStudy (kropot72):